The sum of twice a first number and five times a second number is 111. If the second number is subtracted from five times the first number the result is 21. Find the numbers
Answer by mananth(16949)

The sum of twice a first number and five times a second number is 111.
let one number be x: other number be y
2x+5y=111................ equation 1
..
the second number is subtracted from five times the first number the result is 21.
5x-y=21................... equation 2
multiply equation 2 by 5 and it to equation 1 to eliminate y
25x-5y=105
..
25x-5y+2x+5y=105+111
27x=216
x=8
..
plug the value of x in equation 2
40-y=21
-y=-19
y=19
..
The numbers are 8 & 19